# Robust Hallucination Detection in LLMs via Adaptive Token Selection

### v1

@@ 6 Conclusion
Hallucination detection is essential for the reliable deployment of LLMs. In this paper, We introduce a joint token selection and hallucination detection approach, HaMI, designed to adaptively identify the most likely to be hallucinated tokens, enhancing the robustness of the detection on generation responses of varying lengths and hallucinated entities. Specifically, our approach incorporates a straightforward yet effective multiple instance learning formulation to automatically highlight salient tokens for the training of more accurate hallucination detectors. Additionally, we also explore integrating uncertainty metrics into the original representations to enrich them with more information about truthfulness. Extensive empirical results demonstrate that HaMI substantially outperforms existing methods across diverse QA datasets and LLMs with various characteristics. Our ablation studies offer additional investigations and insights into different aspects of designed modules. While our experiments primarily focus on the QA domain, the principles underlying our method are task-free, suggesting potential applicability to a broad spectrum of other tasks.

### v2

@@ 6 Conclusion
In this paper, we introduce the very first approach that supports joint token selection and hallucination detection, HaMI, enabling adaptive identification of the most hallucinated tokens for learning an optimal hallucination detector. This helps largely enhance the robustness of the detection on generation responses of varying lengths and hallucinated entities. Specifically, HaMI incorporates a straightforward yet effective MIL formulation to automatically highlight salient tokens that are optimal for training the subsequent hallucination detector. Additionally, we also explore integrating uncertainty metrics into the original representations to enrich them with more information about truthfulness. Extensive empirical results demonstrate that HaMI substantially outperforms existing SotA methods across diverse popular datasets and LLMs. Our ablation studies offer important insights into how different designs in HaMI help improve the robustness of the detection. While our experiments primarily focus on the QA domain, the principles underlying our method are task-free, suggesting potential applicability to a broad spectrum of other tasks.
Appendix A Implementation Details
A.1 Setup
For the main results, our hallucination detector $f_{\theta}(\cdot)$ utilises a two-layer MLP with a hidden dimension of 256. The first linear layer is followed by a BatchNorm and ReLU activations and the second linear layer is accompanied by a sigmoid output. As shown in Figure 3b in our paper, our method is relatively insensitive to variations in the number of layers and the hidden dimension size. At the training stage, we use Adam optimiser. For the input representations of our detector, we utilise representations extracted from a single layer and the layer index is determined by the results of the validation set.
We implement our method using PyTorch 2.6.0 [36] and transformers 4.51.3 [43] and conduct all experiments on NVIDIA A100 GPUs. Based on experimental records, the estimated total training and inference time for each dataset under the main experimental settings is as follows: 0.40 GPU-hours for LLaMA-3.1-8B, 0.45 GPU-hours for Mistral-Nemo-Instruct (12B) and 0.65 GPU-hours for LLaMA-3.13-Instruct-70B with 8-bit quantisation. It is important to note that all these results are based on the assumption that the semantic uncertainty score $P_{\text{uncertainty}}^{c}$ is readily available. Otherwise, computing this score involves generating multiple responses and evaluating entailment across them, which introduces nontrivial latency, approximately 7.6 seconds per question for LLaMA-3.1-8B, 8.6 seconds for Mistral-Nemo-Instruct (12B), and 46.1 seconds for LLaMA-3.13-Instruct-70B.
A.2 Prompts for generation and evaluation
All prompts utilised in our experiments refer to those used by Farquhar et al. [16]. For the QA tasks, we prompt selected LLaMA and Mistral models to generate answers without context in a zero-shot manner. Figure 6 presents the used input prompt. To obtain response labels, we prompt GPT-4.1 [1] with OpenAI API to evaluate the quality of the response. The prompts are illustrated in Figure 7. It is observed that GPT-4.1 can make mistakes for the positive samples as checked with the gold answers. Therefore, we ask GPT-4.1 to rejudge samples labelled as positive and discard samples if the result is inconsistent with the first result.
Figure 6: Prompts for QA tasks without context.
Figure 7: GPT-4.1 evaluation prompts
Appendix B Study on $k$ for the top $k$ highest hallucination score selection
As presented in Section 4.1 in our paper, we select token instances with the top $k$ highest hallucination scores as the salient tokens in each sequence. $k$ is dynamically determined by the generated token length $l$ as follows:
$k=\lfloor r_{k}\times l\rfloor+1,$ (11)
where $r_{k}$ is set to 0.1 by default. In this section, we investigate the sensitivity of our method to the choice of $r_{k}$ in the adaptive token selection process. As shown in Figure 8, HaMI exhibits stable performance across a set of $r_{k}$, ranging from 0.00 to 0.20, on all four datasets with LLaMA-3.1-8B and Mistral-Nemo-Instruct (12B). Notably, smaller $r_{k}$ values already yield strong results, which implies that only a few high-salience tokens are sufficient for effective hallucination discrimination.
Figure 8: Study on different choice of $r_{k}$ across four datasets with two LLMs.
Appendix C Limitations and Broader Impacts
Broader impacts.
With the growing deployment of LLMs, ensuring the truthfulness of their outputs has become a critical challenge, especially in high-stakes domains such as finance and healthcare. To enable effective hallucination detection, in this work, we propose a practical and robust approach that leverages internal representations to identify tokens with a high likelihood of being hallucinated and enables the following detection. We hope that our proposed method, HaMI, can contribute to safer and more reliable real-world deployment of LLMs. Although our research focuses on QA tasks, HaMI can be extended to other tasks as well. For example, in the summarisation task, truthfulness can be assessed sentence by sentence. Verifying the correctness of each generated sentence can be reformulated as an entailment query against the source context, allowing hallucination detection to be performed in a QA-like manner. Furthermore, the similarity between a generated sentence and its corresponding context could serve as an uncertainty metric for internal representation enhancement.
Limitations.
Apart from the aforementioned capabilities of HaMI, there remain some concerns. Unlike some black-box uncertainty-based approaches, our method requires access to internal representations, limiting its deployments to open-source LLMs. Moreover, we also explore integrating uncertainty metrics into the original representations to enhance their discriminative capability on correct and incorrect generations. While effective, the integration strategy is relatively simple and we acknowledge that more sophisticated fusion methods could be explored for better detection performance in future research.